# Exploring Young Stellar Variability with Gaia DR3 light curves

### v1

@@ Acknowledgements.
We acknowledge funding from the European Union’s Horizon 2020 research and innovation program (grant agreement No.101004141, NEMESIS). G.M. acknowledges support from the János Bolyai Research Scholarship of the Hungarian Academy of Sciences. We thank Berry Holl for helpful discussions on the use of Gaia DR3 data, Ann-Marie Cody and Laura Venuti for exchanges about the implementation of the Asymmetry and Periodicity metrics. This research was also supported by the International Space Science Institute (ISSI) in Bern, through ISSI International Team project 521 selected in 2021, Revisiting Star Formation in the Era of Big Data (https://teams.issibern.ch/starformation/) This work made use of Astropy (Astropy Collaboration et al., 2013, 2018, 2022), Scikit-learn: Machine Learning in Python (Pedregosa et al., 2011), Pandas (pandas development team, 2020), Numpy (Harris et al., 2020), and Matplotlib (Hunter, 2007). This work has made use of data from the European Space Agency (ESA) mission Gaia (https://www.cosmos.esa.int/gaia), processed by the Gaia Data Processing and Analysis Consortium (DPAC, https://www.cosmos.esa.int/web/gaia/dpac/consortium). Funding for the DPAC has been provided by national institutions, in particular the institutions participating in the Gaia Multilateral Agreement. This research has made use of the SIMBAD database, operated at CDS, Strasbourg, FranceReferences
- Alencar, S. H. P. et al. (2010) Alencar, S. H. P., Teixeira, P. S., Guimarães, M. M., et al. 2010, A&A, 519, A88
- Andre et al. (1993) Andre, P., Ward-Thompson, D., & Barsony, M. 1993, ApJ, 406, 122
- Astropy Collaboration et al. (2022) Astropy Collaboration, Price-Whelan, A. M., Lim, P. L., et al. 2022, ApJ, 935, 167
- Astropy Collaboration et al. (2018) Astropy Collaboration, Price-Whelan, A. M., Sipőcz, B. M., et al. 2018, AJ, 156, 123
- Astropy Collaboration et al. (2013) Astropy Collaboration, Robitaille, T. P., Tollerud, E. J., et al. 2013, A&A, 558, A33
- Auvergne et al. (2009) Auvergne, M., Bodin, P., Boisnard, L., et al. 2009, A&A, 506, 411
- Baluev (2008) Baluev, R. V. 2008, MNRAS, 385, 1279
- Barrado y Navascués & Martín (2003) Barrado y Navascués, D. & Martín, E. L. 2003, AJ, 126, 2997
- Basri et al. (2022) Basri, G., Streichenberger, T., McWard, C., et al. 2022, ApJ, 924, 31
- Bouvier et al. (1999) Bouvier, J., Chelli, A., Allain, S., et al. 1999, A&A, 349, 619
- Bouvier et al. (2013) Bouvier, J., Grankin, K., Ellerbroek, L. E., Bouy, H., & Barrado, D. 2013, A&A, 557, A77
- Bredall et al. (2020) Bredall, J. W., Shappee, B. J., Gaidos, E., et al. 2020, MNRAS, 496, 3257
- Buder et al. (2024) Buder, S., Kos, J., Wang, E. X., et al. 2024, arXiv e-prints, arXiv:2409.19858
- Cabrit et al. (2006) Cabrit, S., Pety, J., Pesenti, N., & Dougados, C. 2006, A&A, 452, 897
- Carpenter et al. (2001) Carpenter, J. M., Hillenbrand, L. A., & Skrutskie, M. F. 2001, AJ, 121, 3160
- Cody & Hillenbrand (2018) Cody, A. M. & Hillenbrand, L. A. 2018, The Astronomical Journal, 156, 71
- Cody et al. (2017) Cody, A. M., Hillenbrand, L. A., David, T. J., et al. 2017, ApJ, 836, 41
- Cody et al. (2022) Cody, A. M., Hillenbrand, L. A., & Rebull, L. M. 2022, The Astronomical Journal, 163, 212
- Cody et al. (2014) Cody, A. M., Stauffer, J., Baglin, A., et al. 2014, AJ, 147, 82
- Creevey et al. (2023) Creevey, O. L., Sordo, R., Pailler, F., et al. 2023, A&A, 674, A26
- De Angeli et al. (2023) De Angeli, F., Weiler, M., Montegriffo, P., et al. 2023, A&A, 674, A2
- Dembinski & et al. (2020) Dembinski, H. & et al., P. O. 2020
- Distefano et al. (2022) Distefano, E., Lanzafame, A. C., Brugaletta, E., et al. 2022, arXiv e-prints, arXiv:2206.05500
- Evans et al. (2023) Evans, D. W., Eyer, L., Busso, G., et al. 2023, A&A, 674, A4
- Eyer et al. (2023) Eyer, L., Audard, M., Holl, B., et al. 2023, A&A, 674, A13
- Eyer & Genton (1999) Eyer, L. & Genton, M. G. 1999, A&AS, 136, 421
- Facchini et al. (2016) Facchini, S., Manara, C. F., Schneider, P. C., et al. 2016, A&A, 596, A38
- Findeisen et al. (2015) Findeisen, K., Cody, A. M., & Hillenbrand, L. 2015, ApJ, 798, 89
- Fischer et al. (2023) Fischer, W. J., Hillenbrand, L. A., Herczeg, G. J., et al. 2023, in Astronomical Society of the Pacific Conference Series, Vol. 534, Protostars and Planets VII, ed. S. Inutsuka, Y. Aikawa, T. Muto, K. Tomida, & M. Tamura, 355
- Gaia Collaboration et al. (2023a) Gaia Collaboration, Bailer-Jones, C. A. L., Teyssier, D., et al. 2023a, A&A, 674, A41
- Gaia Collaboration et al. (2016) Gaia Collaboration, Prusti, T., de Bruijne, J. H. J., et al. 2016, A&A, 595, A1
- Gaia Collaboration et al. (2023b) Gaia Collaboration, Vallenari, A., Brown, A. G. A., et al. 2023b, A&A, 674, A1
- (33) Gaia ESA Archive. , official webpage, https://gea.esac.esa.int/archive/
- Genova et al. (2000) Genova, F., Egret, D., Bienaymé, O., et al. 2000, A&AS, 143, 1
- Gilmore et al. (2022) Gilmore, G., Randich, S., Worley, C. C., et al. 2022, A&A, 666, A120
- Ginsburg et al. (2019) Ginsburg, A., Sipőcz, B. M., Brasseur, C. E., et al. 2019, AJ, 157, 98
- Greene et al. (1994) Greene, T. P., Wilking, B. A., Andre, P., Young, E. T., & Lada, C. J. 1994, ApJ, 434, 614
- Harris et al. (2020) Harris, C. R., Millman, K. J., van der Walt, S. J., et al. 2020, Nature, 585, 357
- Herbst et al. (1994) Herbst, W., Herbst, D. K., Grossman, E. J., & Weinstein, D. 1994, AJ, 108, 1906
- Hillenbrand et al. (2022) Hillenbrand, L. A., Kiker, T. J., Gee, M., et al. 2022, AJ, 163, 263
- Holl et al. (2023) Holl, B., Fabricius, C., Portell, J., et al. 2023, A&A, 674, A25
- Hourihane et al. (2023) Hourihane, A., François, P., Worley, C. C., et al. 2023, A&A, 676, A129
- Hughes et al. (1992) Hughes, P. A., Aller, H. D., & Aller, M. F. 1992, ApJ, 396, 469
- Hunter (2007) Hunter, J. D. 2007, Computing in Science & Engineering, 9, 90
- James & Roos (1975) James, F. & Roos, M. 1975, Comput. Phys. Commun., 10, 343
- Jiang & Hillenbrand (2024) Jiang, S. D. & Hillenbrand, L. A. 2024, AJ, 167, 221
- Joy (1945) Joy, A. H. 1945, ApJ, 102, 168
- Kozłowski (2016) Kozłowski, S. 2016, ApJ, 826, 118
- Kulkarni & Romanova (2008) Kulkarni, A. K. & Romanova, M. M. 2008, MNRAS, 386, 673
- Lada (1987) Lada, C. J. 1987, in Star Forming Regions, ed. M. Peimbert & J. Jugaku, Vol. 115, 1
- Lakeland & Naylor (2022) Lakeland, B. & Naylor, T. 2022, in The 21st Cambridge Workshop on Cool Stars, Stellar Systems, and the Sun, Cambridge Workshop on Cool Stars, Stellar Systems, and the Sun, 111
- Lomb (1976) Lomb, N. R. 1976, Ap&SS, 39, 447
- Marton et al. (2023) Marton, G., Ábrahám, P., Rimoldini, L., et al. 2023, A&A, 674, A21
- Masci et al. (2019) Masci, F. J., Laher, R. R., Rusholme, B., et al. 2019, PASP, 131, 018003
- McGinnis et al. (2015) McGinnis, P. T., Alencar, S. H. P., Guimarães, M. M., et al. 2015, A&A, 577, A11
- Morales-Calderón et al. (2011) Morales-Calderón, M., Stauffer, J. R., Hillenbrand, L. A., et al. 2011, ApJ, 733, 50
- Mowlavi et al. (2023) Mowlavi, N., Holl, B., Lecoeur-Taïbi, I., et al. 2023, A&A, 674, A16
- Ochsenbein et al. (2000) Ochsenbein, F., Bauer, P., & Marcout, J. 2000, A&AS, 143, 23
- Ordenes-Huanca et al. (2024) Ordenes-Huanca, C., Zoccali, M., Bayo, A., et al. 2024, MNRAS, 533, 841
- pandas development team (2020) pandas development team, T. 2020, pandas-dev/pandas: Pandas
- Pecaut & Mamajek (2013) Pecaut, M. J. & Mamajek, E. E. 2013, ApJS, 208, 9
- Pedregosa et al. (2011) Pedregosa, F., Varoquaux, G., Gramfort, A., et al. 2011, Journal of Machine Learning Research, 12, 2825
- Petrov et al. (2015) Petrov, P. P., Gahm, G. F., Djupvik, A. A., et al. 2015, A&A, 577, A73
- Pettersson et al. (2014) Pettersson, B., Armond, T., & Reipurth, B. 2014, A&A, 570, A30
- Randich et al. (2022) Randich, S., Gilmore, G., Magrini, L., et al. 2022, A&A, 666, A121
- Rimoldini et al. (2023) Rimoldini, L., Holl, B., Gavras, P., et al. 2023, A&A, 674, A14
- Rodriguez et al. (2013) Rodriguez, J. E., Pepper, J., Stassun, K. G., et al. 2013, AJ, 146, 112
- Roelens et al. (2017) Roelens, M., Eyer, L., Mowlavi, N., et al. 2017, MNRAS, 472, 3230
- Romanova et al. (2013) Romanova, M. M., Ustyugova, G. V., Koldoba, A. V., & Lovelace, R. V. E. 2013, MNRAS, 430, 699
- Roquette et al. (2020) Roquette, J., Alencar, S. H. P., Bouvier, J., Guarcello, M. G., & Reipurth, B. 2020, A&A, 640, A128
- Roquette et al. (2025) Roquette, J., Audard, M., Hernandez, D., et al. 2025, arXiv e-prints, arXiv:2501.08486
- Saunders et al. (2006) Saunders, E. S., Naylor, T., & Allan, A. 2006, Astronomische Nachrichten, 327, 783
- Savitzky & Golay (1964) Savitzky, A. & Golay, M. J. E. 1964, Analytical Chemistry, 36, 1627
- Scargle (1982) Scargle, J. D. 1982, ApJ, 263, 835
- Simonetti et al. (1985) Simonetti, J. H., Cordes, J. M., & Heeschen, D. S. 1985, ApJ, 296, 46
- Sousa et al. (2016) Sousa, A. P., Alencar, S. H. P., Bouvier, J., et al. 2016, A&A, 586, A47
- Stauffer et al. (2014) Stauffer, J., Cody, A. M., Baglin, A., et al. 2014, AJ, 147, 83
- Stauffer et al. (2016) Stauffer, J., Cody, A. M., Rebull, L., et al. 2016, AJ, 151, 60
- Stelzer et al. (2007) Stelzer, B., Flaccomio, E., Briggs, K., et al. 2007, A&A, 468, 463
- Tofflemire et al. (2017) Tofflemire, B. M., Mathieu, R. D., Ardila, D. R., et al. 2017, ApJ, 835, 8
- VanderPlas (2018) VanderPlas, J. T. 2018, ApJS, 236, 16
- Venuti et al. (2015) Venuti, L., Bouvier, J., Irwin, J., et al. 2015, A&A, 581, A66
- Venuti et al. (2021) Venuti, L., Cody, A. M., Rebull, L. M., et al. 2021, The Astronomical Journal, 162, 101
- Wang et al. (2023) Wang, X.-L., Fang, M., Herczeg, G. J., et al. 2023, Research in Astronomy and Astrophysics, 23, 075015
- Wenger et al. (2000) Wenger, M., Ochsenbein, F., Egret, D., et al. 2000, A&AS, 143, 9
- White & Basri (2003) White, R. J. & Basri, G. 2003, ApJ, 582, 1109
- Wright et al. (2010) Wright, E. L., Eisenhardt, P. R. M., Mainzer, A. K., et al. 2010, AJ, 140, 1868
- Zechmeister & Kürster (2009) Zechmeister, M. & Kürster, M. 2009, A&A, 496, 577
- Zhao et al. (2012) Zhao, G., Zhao, Y.-H., Chu, Y.-Q., Jing, Y.-P., & Deng, L.-C. 2012, Research in Astronomy and Astrophysics, 12, 723
Appendix A Details on light-curve filtering steps
@@ A.1 Bootstrap test for minimum light curve length
We performed bootstrap tests to determine the minimum number of epochs required to ensure statistically robust variability indices. As a baseline, we selected light curves of GAPS sources that contained the highest number of epochs (typically num_selected_g_fov between 70 and 100 observations). To account for variations of signal-to-noise as a function of observed magnitude, we selected 5 such sources within bins with 0.5 mag width for between 6.5 and 20 mag. For each light curve selected, we computed the interquartile range (IQR) as a function of the total number of epochs while randomly removing epochs until only 10 epochs were left. To quantify the change introduced to the IQR by the removal of epochs, we computed a loss function as:
where is the IQR of the original light curve and the IQR at the step of epoch removal. The loss function represents a percentage difference between the original IQR and the IQR when points are removed.
We repeated the process 10 000 times per light curve to minimise biases introduced by specific epochs while randomly removing epochs. An example of the distribution of loss function values as a function of the number of epochs is shown in Fig.26 for the mag bin . Based on this distribution, we inferred the minimum number of epochs needed to keep the relative changes in the IQR under (34.1%) of the original value for 95% of the light curves in that magnitude bin, which in the example shown in Fig.26 corresponded to 31 epochs. Fig. 27 shows the minimum number of epochs as a function of magnitude for all magnitude bins considered. There is a strong peak between 9.5 and 13.5 mags, meaning sources with this magnitude have an IQR changing more intensely, which is likely related to Gaia’s instrumental effects around magnitude 12. To account for this feature, we applied a two-regime condition for the minimum number of observations and required at least 62 epochs for sources with magnitudes between 9.5 and 13.5 and 37 epochs for others.
@@ A.2 Spurious variability identification procedure
As part of GDR3, Holl et al. (2023) provides a series of parameters calculated to facilitate the identification of spurious variability signals introduced by Gaia’s complex scanning law and varying object scanning angles. These parameters were retrieved for our sources from the vari_spurious_signals table through Gaia ESA Archive . In specific, we employed the pre-calculated Spearman correlation between photometric magnitudes and IPD GoF (Image Processing Data Goodness of Fit), spearman_corr_ipd_g_fov, and between magnitudes and their corrected flux excess factor spearman_corr_exf_g_fov. In both cases, a significant correlation suggests that scanning-law effects are likely present in the light curve.
@@
A.2.1 Timescales in excess due to residual spurious variability
In Fig. 29, we show a distribution of the power of the highest peak in the GLS periodogram as a function of the timescale for that peak (derived in Sect. 4.2.1). An excess of sources detected at specific long timescales (e.g., 100, 50 and 25 d) is noticeable in this distribution. Fig. 29 also shows a normalised frequency distribution of GLS timescales for sources affected by spurious variability identified in the previous section. Considering the timescales with at least 10% frequency among sources with spurious variability points to a correspondence of these with the timescales in excess, suggesting a prevalence of sources with spurious variability in our sample. We have further inspected these sources in the context of spearman_corr_ipd_g_fov and spearman_corr_exf_g_fov parameters discussed in the previous section to address if a different strategy could reduce their incidence. For example, the requirement that rather than both, either of these parameters show stronger correlation or using a lower correlation threshold. However, a relevant population of residual spurious variability sources was found even among sources with correlation values close to zero. In particular, we note that the recurrence of residual spurious variability is minimal within the primary (highly) variable sample defined in Sect. 3.5, but accentuated among sources in the secondary sample. Finally, we minimised this residual contamination by utilising the normalised frequency distribution of GLS timescales for sources affected by spurious variability identified in the previous section to mask the most frequent timescales among sources with spurious variability.
A.3 Procedure for outlier removal
Visual inspection of light curves exhibiting the highest peak-to-peak amplitude revealed that many of these sources had their variability and asymmetry analysis dominated by only one or two epochs beyond the level. In the context of Gaia’s sparse light curves, we must address our capacity to judge whether these are spurious outliers or the results of the sparse sampling of a rapidly occurring variability event. For example, short-duration bursting features reported by Cody et al. (2017) typically span 10–20 days and typically show short duty cycles (time spent bursting). At certain latitudes, Gaia’s scanning law can occasionally yield the occurrence of a dozen observed epochs over a couple of days, a phenomenon often referred to as cusps (an example of Gaia DR3 light curve with a cusp is shown in Fig. 8 for the source ). However, if not coincident with cusps, Gaia’s sparse sampling typically includes only one epoch every 10-20 days, henceforth under-resolving and, likely, often missing short-duration bursts.
@@ A.3 Procedure for outlier removal
Appendix B Structure Function’s blind timescales
@@
In this Appendix, we investigate the range of timescales affected by Gaia’s sparse time coverage. To this end, we employed Eq. 3 to calculate structure functions for all 43 840 sources with enough epochs in their light curves according to the procedure discussed in Sect. 3.2. Fig. 31 shows the distribution of calculated -SF() pairs at a population level, where a lower density of points is clearly seen at values around days. The frequency of -SF() pairs occurrence is also shown in the top panel of Fig. 31. We choose a conservative requirement of at least 40% to define a blind-SF range of timescales in the range days. Nevertheless, we note that the issue of underpopulated SF is more critical for the timescale range of days where any given -SF() exists for less than 10% of sources. GDR3 scanning law (Fig. 1) implies that light curves will have varied sampling. Hence, the definition of this blind timescale range does not mean that timescales in this range cannot be measured with the SF method, but rather that SF will less frequently be well-sampled enough to detect such timescales. Based on the identification of this blind range, Sect. 3.2 includes additional processing steps to identify if an aperiodic timescale in this range came from an adequately sampled SF. Finally, we also note that this sampling issue is only expected to affect the SF method with the underlying assumption of an aperiodic SF model. For GLS periodic timescales, different underlying assumptions of a sinusoidal periodic signal apply, and various previous studies based on GDR3 light curves have corroborated its efficiency in detecting timescales in this range (e.g., Distefano et al. 2022).
Appendix C Description of data table

### v2

@@ Acknowledgements.
We acknowledge funding from the European Union’s Horizon 2020 research and innovation program (grant agreement No.101004141, NEMESIS). G.M. acknowledges support from the János Bolyai Research Scholarship of the Hungarian Academy of Sciences. We thank Berry Holl for helpful discussions on the use of Gaia DR3 data, Ann-Marie Cody and Laura Venuti for exchanges about the implementation of the Asymmetry and Periodicity metrics. This research was also supported by the International Space Science Institute (ISSI) in Bern, through ISSI International Team project 521 selected in 2021, Revisiting Star Formation in the Era of Big Data (https://teams.issibern.ch/starformation/) This work made use of Astropy (astropy:2013; astropy:2018; astropy:2022), Scikit-learn: Machine Learning in Python (scikit-learn), Pandas (reback2020pandas), Numpy (harris2020array), and Matplotlib (Hunter:2007). This work has made use of data from the European Space Agency (ESA) mission Gaia (https://www.cosmos.esa.int/gaia), processed by the Gaia Data Processing and Analysis Consortium (DPAC, https://www.cosmos.esa.int/web/gaia/dpac/consortium). Funding for the DPAC has been provided by national institutions, in particular the institutions participating in the Gaia Multilateral Agreement. This research has made use of the SIMBAD database, operated at CDS, Strasbourg, FranceAppendix A Details on light curve filtering steps
@@ A.1 Bootstrap test for minimum light curve length
We performed bootstrap tests to determine the minimum number of epochs required to ensure statistically robust variability indices. As a baseline, we selected light curves of GAPS sources that contained the highest number of epochs (typically num_selected_g_fov between 70 and 100 observations). To account for variations of signal-to-noise as a function of observed magnitude, we selected 5 sources per bin with 0.5 mag width for between 6.5 and 20 mag. For each light curve selected, we computed the interquartile range (IQR) as a function of the number of epochs while randomly removing epochs until 10 epochs were left. To quantify the change introduced to the IQR by the removal of epochs, we computed a loss function as:
where is the IQR of the original light curve and the IQR at the step of epoch removal. The loss function represents a percentage difference between the original IQR and the IQR when points are removed. We repeated the process 10 000 times per light curve to minimise biases introduced by specific epochs while randomly removing epochs. An example of the distribution of loss function values as a function of the number of epochs is shown in Fig.28 for the mag bin m from where we inferred that a minimum of 31 epochs was required to keep the relative changes in the IQR under (34.1%) for 95% of the light curves in that magnitude bin (white cross).
Fig. 29 shows the minimum number of epochs as a function of magnitude for all magnitude bins considered. There is a strong peak between 9.5 and 13.5 mags, meaning sources with this magnitude have an IQR changing more intensely, which is likely related to instrumental effects of Gaia around magnitude 12. To account for this feature, we applied a two-regime condition for the minimum number of observations and required at least 62 epochs for sources with magnitudes between 9.5 and 13.5 and 37 epochs for others.
@@ A.2 Spurious variability identification procedure
As part of GDR3, Holl2023 provides a series of parameters calculated to facilitate the identification of spurious variability signals introduced by the complex scanning law of Gaia and varying object scanning angles. These parameters were retrieved for our sources from the vari_spurious_signals table through GaiaArchive. In specific, we employed the pre-calculated Spearman correlation between photometric magnitudes and IPD GoF (Image Processing Data Goodness of Fit), spearman_corr_ipd_g_fov, and between magnitudes and their corrected flux excess factor spearman_corr_exf_g_fov. In both cases, a significant correlation suggests that scanning-law effects are likely present in the light curve.
@@
A.3 Timescales in excess due to residual spurious variability
In Fig. 31, we show a distribution of the power of the highest peak in the GLS periodogram as a function of the timescale for that peak (derived in Sect. 4.2.1). An excess of sources detected at specific long timescales (e.g. 100, 50 and 25 d) is noticeable in this distribution. Fig. 31 also shows a normalised frequency distribution of GLS timescales for sources affected by spurious variability identified in the previous section. Considering the timescales with at least 10% frequency among sources with spurious variability points to a correspondence of these with the timescales in excess, suggesting a prevalence of sources with spurious variability in our sample. We have further inspected these sources in the context of spearman_corr_ipd_g_fov and spearman_corr_exf_g_fov parameters discussed in the previous section to address if a different strategy could reduce their incidence. For example, the requirement that rather than both, either of these parameters show stronger correlation or using a lower correlation threshold. However, a relevant population of residual spurious variability sources was found even among sources with correlation values close to zero. In particular, we note that the recurrence of residual spurious variability is minimal within the primary (highly) variable sample defined in Sect. 3.5, but accentuated among sources in the secondary sample. Finally, we minimised this residual contamination by utilising the normalised frequency distribution of GLS timescales for sources affected by spurious variability identified in the previous section to mask the most frequent timescales among sources with spurious variability.
A.4 Procedure for outlier removal
Visual inspection of light curves exhibiting the highest peak-to-peak amplitude revealed that many sources had their variability and asymmetry analysis dominated by only one or two epochs beyond the level. In the context of the Gaia sparse light curves, we must address our capacity to judge whether these are spurious outliers or the results of the sparse sampling of a rapidly occurring variability event. For example, short-duration bursting features reported by Cody2017 typically span 10–20 days and typically show short duty cycles (time spent bursting). At certain latitudes, the Gaia scanning law can occasionally yield the occurrence of a dozen observed epochs over a couple of days, a phenomenon often referred to as cusps (see, for example, source in Fig. 9). However, if not coincident with cusps, the sparse sampling of Gaia typically includes only one epoch every 10-20 days, henceforth under-resolving and, likely, often missing short-duration bursts.
@@ A.4 Procedure for outlier removal
Appendix B Blind timescales
@@
In this Appendix, we investigate the range of timescales affected by the sparse time coverage of Gaia. To this end, we employed Eq. 3 to calculate structure functions for all 43 840 sources with enough epochs in their light curves according to the procedure discussed in Sect. 3.2. Fig. 33 shows the distribution of calculated -SF() pairs at a population level, where a lower density of points is clearly seen at values around days. The frequency of -SF() pairs occurrence is also shown in the top panel of Fig. 33. We choose a conservative requirement of at least 40% to define a blind-SF range of timescales in the range days. Nevertheless, we note that the issue of underpopulated SF is more critical for the timescale range of days where any given -SF() exists for less than 10% of sources. GDR3 scanning law (Fig. 1) implies that light curves will have varied sampling. Hence, the definition of this blind timescale range does not mean that timescales in this range cannot be measured with the SF method, but rather that SF will less frequently be well-sampled enough to detect such timescales. Based on the identification of this blind range, Sect. 3.2 includes additional processing steps to identify if an aperiodic timescale in this range came from an adequately sampled SF. Finally, we also note that this sampling issue is only expected to affect the SF method with the underlying assumption of an aperiodic SF model. For GLS periodic timescales, different underlying assumptions of a sinusoidal periodic signal apply, and various previous studies based on GDR3 light curves have corroborated its efficiency in detecting timescales in this range (e.g. Distefano2022).
Appendix C YSO properties
C.1 Ongoing accretion
We used the data described in Sect. 2.2.2 to classify sources as CTTS or WTTS based on their H emission. For equivalent width measurements, we employed spectral type-dependent criteria by BarradoYNacascuesMartin2003 converted into Gaia - colours using the PecautMamajek2013 empirical sequence. For full widths at 10% of the peak flux measurements, we used the WhiteBasri2003 criterion to distinguish between CTTS and WTTS. For 64 sources, we used qualitative flags that report H in emission (from Pettersson2014AandA...570A..30P). When multiple H measurements were available, we considered that one measurement indicating a CTTS classification was enough to grant the source a classification as such. Specifically for EW(H) measurements from ESP-ELS Gaia DR3 we have restricted our analysis to 7 318 sources with the absolute value of reported EW(H) larger than at least 5 times the reported uncertainties. This decision was motivated by the limitations due to / spectral low resolution (; BPRPspectra2023AA...674A...2D) and its representation as Hermite function bases, where the identification of sources with large values of EW(H) are reliable, but where low accreting and non-accreting sources with EW(H) values closer to the stellar continuum may be confused within uncertainties.
C.2 Infrared indices
To estimate we followed a similar procedure as the one in the NEMESIS YSO catalogue (roquette2024): We used the Vizier Photometry viewer tool to retrieve for each source all previously published photometric data available at the Centre de données astronomiques de Strasbourg (Genova2000AAS..143....1G_CDS; Ochsenbein2000AAS..143...23O), and used all data available for the wavelength range m to fit a linear model and measure , with a requirement of at least two photometric data points 2 apart in this interval. Next, we adapted the standard classification scheme for YSOs based on their (Lada:87a; Andre1993; Greene1994) by grouping sources based on the dominant contributor to their SED: sources with significant envelope contribution have (1 378 sources), sources with a thick disc have (6 427 sources), sources with thin discs have (8 743 sources) and discless sources have , (11 631 sources).
C.3 Spectral types
In GDR3, effective temperatures from the general stellar parameterizer from photometry(GSP-Phot; BPRPspectra2023AA...674A..27A) were available for 21 112 of our sources. We combined these with the -Spectral Type scale for PMS stars by PecautMamajek2013 to infer spectral types for our sample. For 7 067 sources, effective temperatures were unavailable, and instead we used a Spectral Type version of the PecautMamajek2013 scale. For this end, individual extinctions were inferred at the position and the GDR3 distance of sources using the Galactic 3D dust map by Edenhofer2024A&A...685A..82E with the dustmaps Python package (DustMaps2018JOSS....3..695M). Edenhofer2024A&A...685A..82E provides a robust map for sources up to 1.25 kpc, with a secondary map based on less data available up to 2 kpc. For 84 sources located at larger distances, we adopted the 2 kpc extinction value as a bottom limit. Although informative, stellar parameters for young stars provided by GSP-Phot have a series of caveats (see BPRPspectra2023AA...674A..27A, for a in-depth discussion), including the use of time-averaged mean BP/RP spectra, which neglects YSO variability, and known degeneracy between effective temperature and extinction - expected to be especially relevant for sources in regions of enhanced line-of-sight extinction such as star-forming regions. Due to these known limitations, we did not attempt to derive sub-spectral types, and we stress that the reported spectral types are rough estimations aimed at a broad insight into the content of our YSO sample.
Appendix D Description of the data table